# Equal contributions

### v2

@@ D D 0(0) 0(0)
We also measured the different combinations of training and inference stages cumulatively, by creating 1,000 models under four conditions. These are deterministic training and deterministic inference, deterministic training and non-deterministic inference, non-deterministic training and deterministic inference, and lastly non-deterministic training and non-deterministic inference. The $V_{\text{ermv}}$ and $V_{c}$ for each of these experiments is presented in Table VI. As expected, the non-deterministic training with non-deterministic inference has the most severe variations, and while training seems to incur more variability, inference variability contributes a non-negligible amount. The runtime for the GraphSAGE model training for the full 10 epochs is $0.48(1)$ seconds with a deterministic operation and $0.18(1)$ seconds with a non-deterministic one. The inference runtimes for a single input are given in Table . Here we find that the deterministic implementation is slower than the non-deterministic one. For this neural network, inference on the LPU accelerator is 30 times faster than the fastest PyTorch implementation, consistent with results presented in [27].
TABLE VII: H100 and Groq kernel runtime for deterministic and non deterministic inference of the GraphSAGE model.
@@ VI Conclusions
The analysis presented here highlights the variability caused by non-deterministic operations for simple parallel reductions and PyTorch operations, and the adverse effects on a full DL training and inference pipeline when using non-deterministic functions. We found that the variability from non-deterministic reductions can approach the tolerance thresholds used in high-accuracy molecular simulation correctness tests, and for DL, will produce a unique and non-reproducible set of model weights with each run using identical inputs. These effects were pronounced for scatter_reduce and index_add operations, which are used extensively in GNNs. GNNs have become a key tool for DL used in molecular simulation. In terms of productivity and performance, using deterministic programming patterns may require more effort or experience, especially to maintain performance, and may not be attainable for all algorithms. Within large DL framework stacks, providing deterministic solutions is left to the developers. We found that documentation on which functions in PyTorch are deterministic may be erroneous or incomplete, highlighting the large burden placed on end users in development of reproducible DL workflows. The benefits of deterministic chip designs, therefore, extend beyond performance improvements, to reproducibility facilitation and support, increasing productivity and correctness. The non-determinism studied here results from asynchronous atomic operations on GPUs, but in HPC and distributed settings there will also be inter-chip and inter-node communication, such as with MPI, leading to more runtime variation. On the LPU architecture, inter-chip communication can be software scheduled, removing such communication variations [28]. Future work should explore effects such as this in addition to the ones presented here.
@@ VII Acknowledgements
This work was supported in part by by the Oak Ridge National Laboratory’s AI LDRD Initiative and used resources of the OLCF, a DOE Office of Science User Facility [DE-AC05-00OR22725].

### v3

@@ D D 0(0) 0(0)
We also measured the different combinations of training and inference stages cumulatively, by creating 1,000 models under four conditions. These are deterministic training and deterministic inference, deterministic training and non-deterministic inference, non-deterministic training and deterministic inference, and lastly non-deterministic training and non-deterministic inference. The $V_{\text{ermv}}$ and $V_{c}$ for each of these experiments is presented in Table VII. As expected, the non-deterministic training with non-deterministic inference has the most severe variations, and while training seems to incur more variability, inference variability contributes a non-negligible amount. The runtime for the GraphSAGE model training for the full 10 epochs is $0.48(1)$ seconds with a deterministic operation and $0.18(1)$ seconds with a non-deterministic one. The inference runtimes for a single input are given in Table VIII; the deterministic implementation is slower than the non-deterministic one on GPU. Inference on the LPU accelerator is 30 times faster than the fastest PyTorch implementation, consistent with results presented in [29].
TABLE VIII: H100 and Groq kernel runtime for deterministic and non deterministic inference of the GraphSAGE model.
@@ VI Conclusions
The analyses presented here highlight variability from non-deterministic kernels for basic parallel reductions, PyTorch functions, and resulting adverse effects on a full DL training and inference pipeline. The variability from non-deterministic reductions can approach the tolerance thresholds used in high-accuracy molecular simulation correctness tests, and for DL, will produce a unique and non-reproducible set of model weights with each run using identical inputs. These effects were pronounced for scatter_reduce and index_add operations, which are used extensively in GNNs. GNNs have become a key tool for DL models used in molecular simulation. In terms of productivity and performance, using deterministic programming patterns may require more effort or experience, especially to maintain performance, and may not be attainable for all algorithms. Within large DL framework stacks, providing deterministic solutions is left to the developers. We found that documentation on which functions in PyTorch are deterministic may be erroneous or incomplete, highlighting the large burden placed on end users in development of reproducible DL workflows. The benefits of deterministic chip designs, therefore, may extend beyond performance improvements, to reproducibility facilitation and support, increasing productivity and correctness. The non-determinism studied here results from asynchronous atomic operations on GPUs, but in HPC and distributed settings there will also be inter-chip and inter-node communication, such as with MPI, leading to more runtime variation. On the LPU architecture, inter-chip communication can be software scheduled, removing such communication variations [30]. Future work should explore effects such as this in addition to the ones presented here. In addition, to mitigate the non-reproducibility produced by the training portions of a DL workflow, deterministic hardware for training could also be proposed; emerging DL chip designs may support such solution and can be evaluated.
@@ VII Acknowledgements
This work was supported in part by the ORNL AI LDRD Initiative and in part by Swiss Platform For Advanced Scientific Computing (PASC), and used resources of the OLCF, a DOE Office of Science User Facility [DE-AC05-00OR22725], and Swiss National Supercomputing Centre (CSCS).